Write the equation x(x - 7) + 10 = 0 in quadratic form and then solve it by factoring.
Answer by checkley71(8403)

X(X-7)+10=0
X^2-7X+10=0
(X-5)(X-2)=0
X-5=0
X=5 ANSWER.
X-2=0
X=2 ANSWER.